Campbell Manufacturing Company (CMC) was started when it acquired $95,000 by issuing common stock. During the first year of operations, the company incurred specifically identifiable product costs (materials, labor, and overhead) amounting to $63,000. CMC also incurred $58,500 of engineering design and planning costs. There was a debate regarding how the design and planning costs should be classified. Advocates of Option 1 believe that the costs should be classified as general, selling, and administrative costs. Advocates of Option 2 believe it is more appropriate to classify the design and planning costs as product costs. During the year, CMC made 4,500 units of product and sold 3,700 units at a price of $38.00 each. All transactions were cash transactions.

Required:

a-1. Prepare an income statement and balance sheet under option 1.

a-2. Prepare an income statement and balance sheet under option 2.

b. Identify the option that results in financial statements that are more likely to leave a favorable impression on investors and creditors.

c. Assume that CMC provides an incentive bonus to the company president equal to 13 percent of net income. Compute the amount of the bonus under each of the two options. Identify the option that provides the president with the higher bonus.

d. Assume a 30 percent income tax rate. Determine the amount of income tax expense under each of the two options. Identify the option that minimizes the amount of the company’s income tax expense.

PLEASE LOOK AT NUMBERS. THE LAST PERSON GAVE ME ALL WRONG NUMBERS. THANK YOU!
Answer #1

Ans) A1 & A2

Income statement Option 1 Option 2
Units sold 3700 3700
Sales (3700*$38) 140600 140600
less;Cost of Goods Sold( Unit product cost * units sold) 51800 99900
Gross Margin 88800 40700
less; Selling & adm Exp 58500
Net Operating Income/Loss 30300 40700

Calculation;

Unit Product cost= Total Production cost/no of produced unit

Option1= 63,000/4500= 14

Option2= 63000+58500/4500=27

Balance Sheet Option 1 Option 2
Assets
Cash (total Liability - Inventory) 114100 114100
Inventory (unsold units * Per unit product cost) 11200 21600
Total Current Assets 125300 135700
Total Assets 125300 135700
Liabilities
Common Stock 95000 95000
Retained Earnings 30300 40700
Total Liability & Equity 125300 135700

b) As the Net Operation Income is high in case of Option B, therefore it will leave a favorable impression on investors and creditors.

c) Option 2 provides higher amount of bonus.

Option 1 Option 2
Net Operating Income 30300 40700
Bonus @ 13% 3939 5291

d) Option 1 minimizes the amount of the company’s income tax expense.

Option 1 Option 2
Net Operating Income 30300 40700
tax @ 30% 9090 12210